OpenStudy (anonymous):

Using a directrix of y = -2 and a focus of (1, 6), what quadratic function is created? f(x) = 1/8 (x - 1)^2 - 2 f(x) = -1/8 (x + 1)^2 - 2 f(x) = -1/16 (x + 1)^2 - 2 f(x) = 1/16 (x - 1)^2 + 2 I think it's the first?

OpenStudy (anonymous):

It is the last one. Actually we can find the equation by equation the distances from a point (x,y) to the focus and to the directirx.

OpenStudy (anonymous):

Let me show you how \[ (y+2)^2=(x-1)^2+(y-6)^2\\ \text { after some simplification you get}\\ -x^2+2 x-33=-16 y\\ y=\frac{1}{16} \left(x^2-2 x+33\right)=\frac{1}{16} \left(x^2-2 x+1 +32\right)=\\ \frac{1}{16} \left((x-1)^2 +32\right)=\frac{1}{16} (x-1)^2 +2 \]

OpenStudy (anonymous):

Thank you very much ! (:

OpenStudy (anonymous):

YW